Use the elimination method to solve the system. 5x – 3y = -22 2 + y = 2
Answer #1

50-39 =-22 (ล X + y = 2 erox1 5x - 34 - - – 22 6 eq @x3 = 3x + 3y = 8x = -16 -16 - 8 N = -2 2 x+y = 2 y = 2-a y=2-(-2) Y=2+2